# Accelerating FRB Search: Dataset and Methods

### v1

@@ 8 Conclusion
The dataset is publicly and freely available to the scientific community (https://www.scidb.cn/en/s/VfaYNr). Based on FAST-FREX, a FRB search algorithm challenge will be held to accelerate the research of machine learning algorithms in this field. Meanwhile, we hope that the FAST-FREX dataset serves not only as a facilitator for AI applications in FRB search but also as a catalyst for the advancement of AI for Science.
This work is partially supported by the National Natural Science Foundation of China (NSFC) (11988101). Di Li is a New Cornerstone investigator. This work is also supported by Key R&D Program of Zhejiang (2024SSYS0012). This work made use of the observation data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences.
@@ Appendix A FAST-FREX Datasheet
In this section, we will comply with the guidelines presented in ”Datasheets for Datasets”  [^6] https://arxiv.org/abs/1803.09010 to provide detailed documentation of the FAST dataset for Fast Radio bursts EXploration (FAST-FREX). Therefore, we will address specific sections that include queries related to the dataset’s motivation, composition, collection process, pre-processing/cleaning/labeling, options for usage, distribution, and maintenance. We will distinguish each section’s primary and subsequent questions using bold and italic fonts.
@@ A.1 Motivation
The Research Center for Astronomical Computing of Zhejiang Laboratory created the dataset. The Zhejiang Laboratory and National Astronomical Observatories, Chinese Academy of Sciences (NAOC) are on behalf of the dataset.
@@
Astronomical Big Data Service Platform Project of the Zhejiang Laboratory funded this dataset. And this work is partially supported by the National Natural Science Foundation of China (NSFC) (11988101). This work is also supported by Key R&D Program of Zhejiang (2024SSYS0012).
@@ A.2 Composition
Each instance includes observation data in FITS  [^7] https://fits.gsfc.nasa.gov/ format with 60 * 1024 time sampling points, a 98.304 $\mu$m or 49.152 $\mu$s sampling rate, and a single polarization channel. The size of each sample file is approximately 244 MB or 488MB, depending on its sampling rate. Parameter files are stored in CSV format to record various FRB parameters of positive sample files.
@@ A.3 Collection Process
The raw observation data were collected from three FRB sources. The FAST continuous monitoring of FRB121102 began in August 2019. During 59.5 hours, 1,652 independent burst events were detected from August 29th to October 29th. Meanwhile, FAST observed FRB20180301 on 2021 March 4th, 9th, and 19th and detected five bright radio bursts. 881 bursts detected from FRB20201124 by FAST between 2021 September 25th and 28th. The data was gathered using 4096 frequency channels over 1.05 GHz to 1.45 GHz, with 0.122 MHz frequency resolution. FRB20121102 has a 98.304 $\mu$s sampling rate, while others have a 49.152 $\mu$s sampling rate. These channels recorded four polarization signals.
@@ A.6 Distribution
The URL to access the dataset is https://www.scidb.cn/en/s/VfaYNr.
@@ A.7 Maintenance
Please reach out to Huaxi Chen (chenhuaxi@zhejianglab.com) and Xuerong Guo (guoxr@zhejianglab.com), the authors of this paper.
@@ C.1 PRESTO
1# RFI Removal
2rfifind FRB20121102_0001.fits -o test1 -time 1
3
4# De-dispersion
5prepsubband -nobary -numout 524288 -nsub 1024 -lodm 350.0 -dmstep 1 -numdms 300 -downsamp 1 -mask test1_rfifind.mask -o FRB20121102_0001 FRB20121102_0001.fits
6
7# Single-pulse Search
8ls *.dat |xargs -n 408 -P 12 python single_pulse_search.py -b -m 2 -t 3.0
@@ C.3 RaSPDAM
The RaSPDAM was implemented in Python. The source code and model file can be found in the attachment. The following example demonstrates an evaluation using a FITS file.
@@
2python3 slope_detection.py -o results FRB20121102_0001.fits

### v2

@@ 8 Conclusion
The dataset is publicly and freely available to the scientific community (https://doi.org/10.57760/sciencedb.15070 (catalog DOI: 10.57760/sciencedb.15070)). Based on FAST-FREX, a FRB search algorithm challenge will be held to accelerate the research of machine learning algorithms in this field. Meanwhile, we hope that the FAST-FREX dataset serves not only as a facilitator for AI applications in FRB search but also as a catalyst for the advancement of AI for Science.
This work is partially supported by the National Natural Science Foundation of China (NSFC) (12588202). This work is also supported by Key R&D Program of Zhejiang (2024SSYS0012) and the National Key R&D Program of China(2022YFB4501405). This work made use of the observation data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences.
@@ Appendix A FAST-FREX Datasheet
In this section, we will comply with the guidelines presented in ”Datasheets for Datasets” (Gebru et al., 2018) [^6] https://arxiv.org/abs/1803.09010 to provide detailed documentation of the FAST dataset for Fast Radio bursts EXploration (FAST-FREX). Therefore, we will address specific sections that include queries related to the dataset’s motivation, composition, collection process, pre-processing/cleaning/labeling, options for usage, distribution, and maintenance. We will distinguish each section’s primary and subsequent questions using bold and italic fonts.
@@ A.1 Motivation
The Research Center for Astronomical Computing of Zhejiang Laboratory created the dataset. The Zhejiang Laboratory and National Astronomical Observatories, Chinese Academy of Sciences are on behalf of the dataset.
@@
This work is partially supported by the National Natural Science Foundation of China (NSFC) (12588202). This work is also supported by Key R&D Program of Zhejiang (2024SSYS0012) and the National Key R&D Program of China(2022YFB4501405).
@@ A.2 Composition
Each instance includes observation data in FITS  [^7] https://fits.gsfc.nasa.gov/ format with 60 * 1024 time sampling points, a 98.304 $\mu$m or 49.152 $\mu$s sampling rate, and a single polarization channel. Parameter files are stored in CSV format to record various FRB parameters of positive sample files.
@@ A.3 Collection Process
The raw observation data were collected from three FRB sources by FAST. The data was gathered using 4096 frequency channels over 1.05 GHz to 1.45 GHz, with 0.122 MHz frequency resolution. FRB20121102 has a 98.304 $\mu$s sampling rate, while others have a 49.152 $\mu$s sampling rate. These channels recorded four polarization signals.
@@ A.6 Distribution
The dataset can be accessed viahttps://doi.org/10.57760/sciencedb.15070 (catalog DOI: 10.57760/sciencedb.15070).
@@ A.7 Maintenance
Please reach out to Huaxi Chen (chenhuaxi@zhejianglab.org) and Xuerong Guo (guoxr@zhejianglab.org), the authors of this paper.
@@ C.1 PRESTO
1 # RFI Removal
2 rfifind FRB20121102_0001.fits -o test1 -time 1
1 # De-dispersion
2 prepsubband -nobary -numout 524288 -nsub 1024 -lodm 350.0 -dmstep 1 -numdms 300 -downsamp 1 -mask test1_rfifind.mask -o FRB20121102_0001 FRB20121102_0001.fits
1 # Single-pulse Search
2 ls *.dat |xargs -n 408 -P 12 python single_pulse_search.py -b -m 2 -t 3.0
@@ C.3 RaSPDAM
The RaSPDAM was implemented in Python. The source code is released on https://github.com/zhejianglab/RaSPDAM. The following example demonstrates an evaluation using a FITS file.
@@
2python3 raspdam.py FRB20121102_0001.fits